OpenStudy (anonymous):

Sally is building a flower bed in the shape of a 30°-60°-90° triangle. The hypotenuse is 18 feet. How long is the longest leg? Sally is building a flower bed in the shape of a 30°-60°-90° triangle. The hypotenuse is 18 feet. How long is the longest leg? @Mathematics

OpenStudy (diyadiya):

longest leg would be the side opposite the 60 degree angle, sine(60) = x/18 x = 18 * sine(60) = 18 sqrt3/2 = 9 sqrt3

OpenStudy (anonymous):

The given triangle is a right angled triangle and In a right angled triangle, the hypotenuse is the longest side. So, the longest leg is 18 ft..... (hypotenuse is given as 18 ft..)

OpenStudy (anonymous):

\[18\text{Cos}[30{}^{\circ}]=9 \sqrt{3}=15.5885 \]

OpenStudy (anonymous):

We need the "longest" side.... How can 15.5885 be longer than 18 ????????

OpenStudy (anonymous):

15.5 feet is less than the hypotenuse which is given as 18 feet.

OpenStudy (anonymous):

That is what I am saying... The question asks "How long is the longest leg? " so the correct answer is 18 ft.

OpenStudy (anonymous):

Don't confuse "longest leg" with "longest side" The legs of a right triangle are the sides left when the hypotenuse is removed.